OpenStudy (anonymous):

y = 4 3 x + 3 y = − 2 3 x − 3

OpenStudy (anonymous):

I am not sure what y = − 2 3 x − 3 means?

OpenStudy (anonymous):

4 over 3times x plus 3 and negative 2 thirds times x -3

OpenStudy (anonymous):

systems of quations

OpenStudy (anonymous):

equations*

OpenStudy (anonymous):

y = (4/3)x + 3 y = (-2/3)x - 3 Let's use the subtraction method to get 0 = (6/3)x + 6 Okay with this?

OpenStudy (anonymous):

subtraction method or substitution method?

OpenStudy (anonymous):

Let's use the subtraction method to get 0 = (6/3)x + 6 Okay with this?

OpenStudy (anonymous):

x=9?

OpenStudy (anonymous):

Hmmmm ...... 0 = (6/3)x + 6 six divided by 3 is 2 0 = 2x +6 subtract 6 from both sides -6 = 2x x = ??

OpenStudy (anonymous):

y=6/3x+6 i got that far then stumble

OpenStudy (anonymous):

Did you see my last post? I worked most of it out for you ... 0 = (6/3)x + 6 six divided by 3 is 2 0 = 2x +6 subtract 6 from both sides -6 = 2x x = ??

OpenStudy (anonymous):

i can't figure out x im getting like

OpenStudy (anonymous):

so y = -3?

OpenStudy (anonymous):

i meen x = -3